# SoK: Speedy Secure Finality

### v1

@@ 3.3 Gasper: Ethereum’s Ebb-n-Flow Protocol
Definition 3.5 (Epoch Boundary Pairs).
@@ 3.5 Attacks on Gasper
Balancing attacks have been discovered [attacks] that can allow a validator controlling a small fraction of total network stake and no control over network delay to achieve long-range reorg attacks.
@@
4 Path towards Single-Slot Finality
@@
4.1 Goldfish: A drop in replacement for LMD-GHOST
The balancing and ex-ante attacks possible on Gasper are primarily because not all attesters share the same view as the proposer at the time of the attestation. To address these issues, a new dynamically available protocol was proposed by D’Amato et al. called Goldfish [goldfish].
@@ 4.1 Goldfish: A drop in replacement for LMD-GHOST
Definition 4.1 (Message Buffering).
@@
Definition 4.2 (Vote Expiry).
@@
4.2 RLMD-GHOST: Providing Asynchrony Resilience
@@
To address the challenges of asynchrony resilience, D’Amato and Zanolini introduced RLMD-GHOST [rlmd-ghost] which is a dynamically available protocol with asynchrony resilience. Note that challenges to asynchrony arise only because of the strict vote expiry proposed in Goldfish. RLMD-GHOST relaxes the vote expiry to balance dynamic availability with asynchrony resilience.
@@
D’Amato and Zanolini also specify the joining protocol for the propose-vote-merge family [rlmd-ghost]. When a sleepy validator $V$ wakes up in slot $t$, it receives all the messages while it was asleep and adds them to its buffer $\mathcal{B}_{V}$. It then waits for the next merge phase $3t\Delta+2\Delta$ to merge its $\mathcal{B}_{V}$ into its view $\mathcal{T}_{V}$. Until this point, the validator is considered dreamy [goldfish] and only after the merge is it allowed to participate in the protocol as a regular awake node.
@@
4.3 Generalized Sleepy Model
RLMD-GHOST is analyzed in the generalized sleepy model introduced by D’Amato and Zanolini [rlmd-ghost], which refines the original Pass–Shi sleepy model [sleepy] by accounting for the age of votes that may be safely incorporated into a validator’s view. This refinement allows us to quantify precisely how the vote-expiry parameter $\eta$ influences both dynamic availability and asynchrony resilience.
@@ 4.3 Generalized Sleepy Model
The original sleepy model [sleepy] requires only that the majority of online validators remain honest in each slot. The generalized model quantifies how much “stale” information can be safely incorporated into the fork-choice rule via $\eta$.
Definition 4.3 (Degree of Dynamic Availability).
@@
4.4 Single-Slot Finality on Ethereum
Gasper still takes between 64 and 95 slots to finalize blocks. This is a huge concern for several applications which require faster confirmations. Building on top of RLMD-GHOST, D’Amato and Zanolini proposed a simple single-slot finality protocol [ssf] which combines the synchronous dynamically available protocol with a partially synchronous finality gadget.
@@ 4.4 Single-Slot Finality on Ethereum
5 Endgame for Fast Finality
@@
As Vitalik points out in his blog, There is a deep philosophical reason why epoch-and-slot architectures seem to be so hard to avoid: it inherently takes less time to come to approximate agreement on something, than to come to maximally-hardened ”economic finality” agreement on it. [vitalik]
@@
Reducing slot times also falls under this goal. One of the most well understood protocols that aim at providing fast finality while also reducing slot times is 3-Slot Finality protocol or 3SF proposed by D’Amato et al. [3sf] The original SSF protocol had two voting phases per slot while 3SF proposes just one voting phase per slot while maintaining all safety and liveness guarantees of SSF.
5.1 3-Slot Finality
The 3SF protocol merges the head and FFG votes into a single phase in an attempt to reduce slot times. This also means that finality is achieved in 3 slots instead of the same slot. This protocol is also divided into 4 phases:
@@
    PROPOSE ($4t\Delta$): In this phase, the block proposer uses RLMDGHOSTRLMD-GHOST similarly to SSF and freezes its view with a small difference that vote expiry applies to the combined head and FFG vote described in the next phase.
@@
    VOTE ($4t\Delta+\Delta$): In this phase, the validators send the vote similar to in SSF but also send an FFG-VOTE from the greatest justified checkpoint to the tip of the chain calculated by the $RLMD$ rule.
@@
    CONFIRM ($4t\Delta+2\Delta$): In this phase, validators check if at least a supermajority can cast the VOTE message. If yes, it updates the available chain $chAva$ with the confirmed chain $chConf$ if $chConf$ is a prefix of $chAva$.
@@
    MERGE ($4t\Delta+3\Delta$): In this phase, all validators merge their frozen views to be used in the following slot.
A block $B_{t}$ is proposed in slot $t$, and block $B_{t+1}$ is proposed in slot $t+1$ which is a direct child of $B_{t}$. Validators voting in slot $t+1$ see $B_{t}$ as justified if a supermajority link is established via fast confirmations on $B_{t}$. Then in the next slot $t+2$, if $B_{t+1}$ is seen as justified by validators in $t+2$ and they submit a supermajority link from $B_{t+1}$ to $B_{t+2}$, $B_{t}$ gets finalized.
Finally, we must note that Gasper or SSF used FFG-votes and slashing conditions as described in Casper. This is sufficient for these protocols as it is not possible for two checkpoints from the same slot to be justified. However, this is not true for 3SF. Therefore, in order to keep the protocol safe, an additional slashing condition must be included:
  • validator in the network is said to violate the protocol if it makes two distinct attestations corresponding to two links $s_{1}\rightarrow t_{1}$ and $s_{2}\rightarrow t_{2}$ such that $slot(s_{1})<slot(s_{2})$ and $h_{t_{2}}<h_{t_{1}}$.
With this, 3SF protocol achieves the same safety and liveness guarantees as Gasper. Note that 3SF is also a protocol without any subsampling which means that its implementation is also dependent on reducing the number of signatures required to aggregate in a slot.
5.2 Bottleneck of Signatures
As mentioned previously, after a certain number of validators in the network, the aggregation of signatures becomes the bottleneck in slot times. When Gasper was being developed for the Merge, slot and epoch based architecture was used solely because it was not possible to aggregate so many signatures in a 12 second slot. At the time of the merge, there were around 400,000 validator keys, which shot up to and plateaued around 1 million by 2024.
Breaking an epoch into 32 slots meant each committee only has around 30,000 validators, whose signatures could be collected and aggregated within the slot duration. Any faster finality protocol (especially inspired by RLMD) like 3SF cannot have such subsampling. Currently Ethereum has more than 1 million validators which will not be feasible with any such architecture.
Decentralization, or the number of unique independent nodes in the network, is a very important property and aim for Ethereum. Reducing the number of validator keys may feel like it goes against this idea but in reality that’s not true. Even though Ethereum has more than 1 million validators, the actual number of nodes that run the network is around 8,000 - 10,000. This is because of the maximum effective balance.
Definition 5.1 (Maximum Effective Balance).
The maximum effective balance is the maximum amount of weight allowed for any validator for their attestation. If their staked value is greater than or equal to this value, then the weight of their vote would not be greater than this value.
This value was set to 32 ETH in the Merge. Because of this, entities that wished to stake larger ETH amounts did so by making several validator keys that acted like different validators but are part of the same physical node and controlled by the same entity. This means that if a single entity wanted to deploy 1024 ETH, they can do so by creating 32 validator keys and running all of them under the same node.
This beats the point of independent validators and puts a lot of strain on the network to collect, distribute and aggregate signatures. Estimates that put the number of independent nodes at around 10,000 means there are 100 times more validator keys than nodes because of the maximum effective balance.
@@ 5.2 Bottleneck of Signatures
To prevent this, the limit was recently increased from 32 ETH to 2048 ETH in the Pectra fork as described in EIP-7251 [maxB]. Notice that doing so will allow the validator keys to fall drastically and be closer to the number of real nodes in network today. Although doing so would be a practical challenge because of engineering and maintenance overhead faced by staker networks who manage these validator keys. Achieving this would also make the implementation of a fast finality protocol like 3SF feasible in the Ethereum network.
6 Conclusion

### v2

@@ III-C Gasper: Ethereum’s Ebb-n-Flow Protocol
Definition III.5 (Epoch Boundary Pairs).
@@ III-E Attacks on Gasper
Balancing attacks have been discovered [27] that can allow a validator controlling a small fraction of total network stake and no control over network delay to achieve long-range reorg attacks.
@@
IV Path towards Single-Slot Finality
@@
IV-A Goldfish: A drop in replacement for LMD-GHOST
The balancing and ex-ante attacks possible on Gasper are primarily because not all attesters share the same view as the proposer at the time of the attestation. To address these issues, a new dynamically available protocol was proposed by D’Amato, Nue, Tas and Tse called Goldfish [10].
@@ IV-A Goldfish: A drop in replacement for LMD-GHOST
Definition IV.1 (Message Buffering).
@@
Definition IV.2 (Vote Expiry).
@@
IV-B RLMD-GHOST: Providing Asynchrony Resilience
@@
To address the challenges of asynchrony resilience, D’Amato and Zanolini introduced RLMD-GHOST [15] which is a dynamically available protocol with asynchrony resilience. Note that challenges to asynchrony arise only because of the strict vote expiry proposed in Goldfish. RLMD-GHOST relaxes the vote expiry to balance dynamic availability with asynchrony resilience.
@@
D’Amato and Zanolini also specify the joining protocol for the propose-vote-merge family [15]. When a sleepy validator $V$ wakes up in slot $t$, it receives all the messages while it was asleep and adds them to its buffer $\mathcal{B}_{V}$. It then waits for the next merge phase $3t\Delta+2\Delta$ to merge its $\mathcal{B}_{V}$ into its view $\mathcal{T}_{V}$. Until this point, the validator is considered dreamy [10] and only after the merge is it allowed to participate in the protocol as a regular awake node.
@@
IV-C Generalized Sleepy Model
RLMD-GHOST is analyzed in the generalized sleepy model introduced by D’Amato and Zanolini [15], which refines the original Pass–Shi sleepy model [25] by accounting for the age of votes that may be safely incorporated into a validator’s view. This refinement allows us to quantify precisely how the vote-expiry parameter $\eta$ influences both dynamic availability and asynchrony resilience.
@@ IV-C Generalized Sleepy Model
The original sleepy model [25] requires only that the majority of online validators remain honest in each slot. The generalized model quantifies how much “stale” information can be safely incorporated into the fork-choice rule via $\eta$.
Definition IV.3 (Degree of Dynamic Availability).
@@
IV-D Single-Slot Finality on Ethereum
Gasper still takes between 64 and 95 slots to finalize blocks. This is a huge concern for several applications which require faster confirmations. Building on top of RLMD-GHOST, D’Amato and Zanolini proposed a simple single-slot finality protocol [14] which combines the synchronous dynamically available protocol with a partially synchronous finality gadget.
@@ IV-D Single-Slot Finality on Ethereum
V Endgame for Fast Finality
@@
As Vitalik points out in his blog [6],
“There is a deep philosophical reason why epoch-and-slot architectures seem to be so hard to avoid: it inherently takes less time to come to approximate agreement on something, than to come to maximally-hardened ”economic finality” agreement on it.”
@@
Reducing slot times also falls under this goal. One of the most well understood protocols that aim at providing fast finality while also reducing slot times is 3-Slot Finality protocol or 3SF proposed by D’Amato, Saltini, Tran and Zanolini. [12] The original SSF protocol had two voting phases per slot while 3SF proposes just one voting phase per slot while maintaining all safety and liveness guarantees of SSF.
To clarify the structural commonality between recent fast-finality designs, we present a high-level propose-vote-merge abstraction. This template is not intended as a protocol specification, but as a normal form capturing the minimal control parameters that differentiate existing constructions.
Propose–Vote–Merge Template for Speedy Secure Finality State: local view $\mathcal{T}$, message buffer $\mathcal{B}$ Parameters: vote window $\eta\in\mathbb{N}\cup\{\infty\}$, voting rounds $r\geq 1$ PROPOSE: Merge buffered messages: $\mathcal{T}\leftarrow\mathsf{Merge}(\mathcal{T},\mathcal{B})$ Compute head: $B\leftarrow\mathsf{ForkChoice}(\mathcal{T},\eta)$ Broadcast $\mathsf{Propose}(B,\mathcal{T})$ VOTE ($r$ rounds): Collect votes from slots $t-\eta,\,t-1]$ Update head / justification state Broadcast vote message MERGE: Finalize view $\mathcal{T}$ Discard expired messages from $\mathcal{B}$
Figure 5: A unifying propose–vote–merge abstraction underlying modern fast-finality protocols.
Within this abstraction, Gasper corresponds to an epoch-subsampled instantiation with effectively unbounded vote expiry. Single-slot finality corresponds to $r=2$ with bounded $\eta$, while 3-slot finality collapses the voting phases to $r=1$. This perspective makes explicit that 3SF reduces per-slot communication without introducing new consensus primitives.
V-A 3-Slot Finality (3SF)
The 3SF protocol represents a deliberate retreat from the objective of single-slot finality in favor of a design that is compatible with the dominant engineering bottleneck of large-scale proof-of-stake systems: signature aggregation latency. Rather than optimizing for the minimum theoretical finalization depth, 3SF optimizes for slot length under realistic communication assumptions, while preserving the accountable safety and liveness guarantees of SSF.
At a high level, 3SF merges the head vote and FFG vote phases of SSF into a single voting phase per slot. This reduction removes one full signature-aggregation round from the critical path of each slot. The resulting protocol finalizes blocks deterministically within three consecutive slots under synchrony and sufficient honest participation.
In SSF, a slot contains two voting phases, each of which under aggregation requires approximately $2\Delta$ time. Consequently, the effective slot length is dominated by voting rather than proposal or merge phases. 3SF observes that finality latency and slot latency are distinct optimization targets: sacrificing same-slot finality allows one to shorten the slot itself, which in turn reduces MEV extraction windows and improves throughput-sensitive applications such as AMMs.
Time is divided into slots indexed by $t$, each of duration $4\Delta$, subdivided into four phases:
@@
    $\mathsf{PROPOSE}([4t\Delta,\,4t\Delta+\Delta)$). The slot proposer computes the available chain head using a dynamically available fork-choice rule like RLMD-GHOST [^1] The original 3SF paper described two protocols, one which used RLMD-GHOST as it’s fork-choice rule and another using a quorum based dynamically available protocol called TOB-SVD[13]. The quorum based variant was later spun out from 3SF with some algorithmic tweaks as an independent protocol called Majorum[11]. applied to its frozen view, and broadcasts a proposal extending that chain. As in SSF, vote expiry applies to the combined vote described below.
@@
    $\mathsf{VOTE}([4t\Delta+\Delta,\,4t\Delta+2\Delta)$). Each active validator emits a single vote message that simultaneously serves two roles:
    • a head vote for the dynamically available chain, and
    • an FFG-style vote linking the greatest justified checkpoint in its view to the voted head.
    Formally, a validator broadcasts
    $\textsf{vote}(ch,\,C_{s}\rightarrow C_{t}),$
    where $C_{s}$ is the greatest justified checkpoint and $C_{t}=(ch,t)$ is a target checkpoint for the current slot.
@@
    $\mathsf{CONFIRM}([4t\Delta+2\Delta,\,4t\Delta+3\Delta)$). Validators check whether a supermajority of votes supports a consistent chain prefix. If so, the locally confirmed chain is updated, provided it is a prefix of the available chain.
@@
    $\mathsf{MERGE}([4t\Delta+3\Delta,\,4(t+1)\Delta)$). Buffered messages are merged into the local view, finalizing the state for the next slot.
Let $B_{t}$ denote the block proposed in slot $t$.
  • In slot $t+1$, if a supermajority link $B_{t}\rightarrow B_{t+1}$ is observed, $B_{t}$ becomes justified.
  • In slot $t+2$, if $B_{t+1}$ is justified and a supermajority link $B_{t+1}\rightarrow B_{t+2}$ is formed, then $B_{t}$ is finalized.
Thus, blocks proposed by honest validators are finalized within three slots, independent of proposer honesty in intermediate slots.
Figure 6: Four phases of slot $t$ with 3SF
A subtle but fundamental difference between 3SF and Gasper or SSF arises from the merged vote structure. In Casper FFG and SSF, it is impossible for two distinct checkpoints from the same slot to be justified. In 3SF, this invariant no longer holds: validators may justify multiple checkpoints per slot unless constrained.
To restore accountable safety, 3SF introduces an additional monotonicity condition.
Definition V.1 (3SF Monotonicity Slashing Condition).
A validator is slashable if it produces two votes corresponding to links
$s_{1}\rightarrow t_{1}\quad\text{and}\quad s_{2}\rightarrow t_{2}$
such that
  1. 1.
    $\mathrm{slot}(s_{1})<\mathrm{slot}(s_{2})$, and
  2. 2.
    $\mathrm{height}(t_{2})<\mathrm{height}(t_{1})$.
Intuitively, this condition enforces that a validator’s justified targets must advance monotonically with time. This rule has no analogue in Casper because Casper’s voting structure makes such regressions impossible by construction.
With this additional constraint, 3SF achieves the same accountable safety and plausible liveness guarantees as SSF and Gasper under the standard $f<\tfrac{n}{3}$ fault assumption. Conflicting finalized chains imply slashing of at least one-third of total stake, while liveness holds during sufficiently long synchronous periods with adequate honest participation.
A key conceptual insight exposed by 3SF is that finality depth is not the sole driver of economic security. Reducing slot duration can shrink MEV extraction windows even when finality itself is delayed by additional slots. In this sense, 3SF occupies a distinct point in the design space: it trades confirmation depth for temporal density of blocks.
Moreover, 3SF demonstrates that finality gadgets need not be optimized solely for worst-case latency. Instead, they can be co-designed with aggregation and networking realities, yielding protocols that are asymptotically less elegant than SSF but strictly superior under real-world constraints.
Viewed holistically, SSF, two-slot variants, and 3SF can be compared as follows:
  • SSF minimizes finality depth but maximizes per-slot communication.
  • 3SF minimizes per-slot communication but increases finality depth.
  • Intermediate designs interpolate between these extremes.
V-B Signature Aggregation as a Fundamental Bottleneck
We formalize the engineering limitation imposed by large validator sets and its implications for fast-finality protocols. We present a partial and intentionally abstract formalization whose role is heuristic rather than complete. It is used to guide intuition and structure the discussion of security and performance properties, without claiming coverage of all protocol variants or edge cases, particularly those arising in future 3SF-derived designs.
Definition V.2 (Validator Keys and Independent Nodes).
Let $k$ denote the number of active validator keys participating in consensus and $n$ denote the number of independently operated consensus nodes (i.e., independent failure domains) [^2] Not to be confused with our initial definition of $n$ as total number of consensus nodes. In Ethereum, $k\gg n$ due to stake-splitting across multiple validator keys operated by the same entity.
Definition V.3 (Maximum Effective Balance).
The maximum effective balance $B_{\max}$ is an upper bound on the voting weight of a single validator key. Any stake in excess of $B_{\max}$ must be represented via additional validator keys.
At the Merge, Ethereum set $B_{\max}=32$ ETH, implying that an entity staking $S$ ETH controls at least $\lceil S/32\rceil$ validator keys, regardless of whether these keys are operated on a single physical node.
Theorem V.1 (Signature Aggregation Lower Bound).
In any slot-based proof-of-stake protocol where all validator keys vote in every slot, the time duration of a slot is lower-bounded by $\Omega(k)$ cryptographic signature aggregation.
Proof.
Each validator key produces a distinct signature that must be (i) received, (ii) verified or aggregated, and (iii) redistributed to other participants. Even under optimistic assumptions i.e perfect aggregation trees, no adversarial delays, and constant-size aggregated signatures, the aggregation process must incorporate information from each of the $k$ keys. Therefore, the total aggregation work per slot grows at least linearly in $k$, establishing a lower bound of $\Omega(k)$. ∎
Theorem V.2 provides an empirical proof that fast finality protocols today are infeasible for large number of validators in Ethereum. Assume each validator has bounded inbound bandwidth $W$ (bits/sec) and bounded per slot CPU budget; i.e., there exists a constant $W$ such that no honest node can receive more than $W\tau$ bits within a slot.
Theorem V.2 (Infeasibility of Global-Vote Fast Finality under Bounded Bandwidth).
Let $k$ be the number of active validator keys and let $\tau>0$ be the fixed slot duration. Consider any slot-based proof-of-stake protocol in which a block can be confirmed or finalized in a slot only if an aggregate attestation incorporating votes from at least $qk$ distinct validator keys is produced in that slot, for some constant $q>0$.
Assume that each honest participant has bounded inbound bandwidth $W$ (bits per second). If
$qkb\;>\;W\tau,$
where $b>0$ is a lower bound on the number of bits required per vote contribution, then such a protocol is infeasible. In particular, for $k\approx 10^{6}$ and $\tau=12$ seconds, global-vote fast-finality protocols without subsampling (including SSF- and 3SF-style constructions) are infeasible under current slot-time constraints.
Proof.
For a block to be confirmed in a given slot, some honest participant must be able to determine that at least $qk$ distinct validator keys contributed valid votes in that slot. This requires receiving information sufficient to account for each of these $qk$ contributions.
Let $b>0$ be a lower bound on the number of bits needed to convey one such contribution (e.g., a signature share plus identity or anti-equivocation metadata). Then any participant that verifies the aggregate must receive at least qkbqkb bits during the slot.
Under the bounded-bandwidth assumption, no participant can receive more than $W\tau$ bits in a slot. If $qkb>W\tau$, this necessary condition cannot be met, and confirmation within the slot is impossible. The claim follows by direct contradiction. ∎
Theorem V.3 (Key Inflation without Decentralization Gain).
Under a fixed maximum effective balance $B_{\max}$, increasing total staked ETH inflates the number of validator keys $k$ without increasing the number of independent consensus nodes $n$.
Proof.
Each independent operator with stake $S$ must instantiate at least $\lceil S/B_{\max}\rceil$ validator keys. Increasing $S$ increases $k$ proportionally while leaving the operator count and thus the number of independent failure domains, unchanged. Therefore, $k$ grows independently of $n$. ∎
@@ V-B Signature Aggregation as a Fundamental Bottleneck
One of the ways to enable fast finality in this setup would be to increase the effective balance $B_{max}$ for validators. Notice that doing so will allow the validator keys to fall drastically and be closer to the number of real nodes in network today. Although doing so would be a practical challenge because of engineering and maintenance overhead faced by staker networks who manage these validator keys. Achieving this would also make the implementation of a fast finality protocol like 3SF feasible in the Ethereum network.
Theorem V.4 (Enabling Fast Finality via Effective Balance Scaling).
Increasing the maximum effective balance $B_{\max}$ reduces the per slot signature aggregation burden without weakening fault tolerance.
Proof.
Consider an operator with total stake $S>0$. Under a maximum effective balance $B_{\max}$, the operator must instantiate at least
$k(S,B_{\max}):=\left\lceil\frac{S}{B_{\max}}\right\rceil$
validator keys. The function $k(S,B_{\max})$ is weakly decreasing in $B_{\max}$ for fixed $S$. Summing over all operators, increasing $B_{\max}$ induces a weakly decreasing total number of validator keys $k$ in the system.
The value of $n$ depends only on the partition of total stake across operators and is invariant under changes to $B_{\max}$. Hence, scaling $B_{\max}$ reduces $k$ while leaving $n$ unchanged.
Let $\alpha\in[0,1]$ denote the fraction of total stake controlled by adversarial operators. Safety and liveness guarantees of proof-of-stake consensus depend on $\alpha$ relative to fixed thresholds (e.g., $\alpha<1/3$), not on the number of validator keys used to represent that stake. Since increasing $B_{\max}$ leaves the stake distribution, and thus $\alpha$,unchanged, the fault-tolerance threshold is preserved.
By Theorem V.2, the per slot signature aggregation requirement scales at least linearly in $k$. Therefore, reducing $k$ via an increase in $B_{\max}$ strictly relaxes the aggregation constraint, enabling the feasibility of fast-finality protocols such as 3SF once $k$ is reduced to the same order of magnitude as $n$.
EIP-7251[23] raises $B_{\max}$ from 32 ETH to 2048 ETH, permitting a reduction of $K$ by up to a factor of 64 in the steady state. This shift moves Ethereum from a committee-mandated architecture toward one where global-vote fast-finality protocols are practically viable.
VI Conclusion